It takes Nina 3 hours to proof a chapter of Hawkes Learning Systems' Statistics book and it takes Cynthia 12 hours. How long would it take them together?
Answer by ewatrrr(24785)

 
Hi,
It takes Nina 3 hours to proof a chapter of Hawkes Learning Systems' Statistics
book and it takes Cynthia 12 hours.
Let x represent the time it would take them together
Note: job / hour is the key
1/3 + 1/12 = 1/x |Multiplying thru by 12x so as all denominators = 1
4x + x = 12
5x = 12
x = 2.4 hrs, the time it would take them together